Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
gherardini - 16:01 Wednesday 16 June 2021 (52164) Print this report
Operator Report - Afternoon shift

This morning was dedicated to ISC activity; Manuel and Michele worked all the shift without any major problem, the ISC work will go on in the afternoon...


other activities of this morning:

- SQZ: FCEM baffle installation and filter cavity alignment recovery;
- EDB cabling from 7:30UTC to 10:00UTC;
- ENV: west end environmental sensor installation, done at around 10:30UTC;

SUSP
one of the MC F0 coil is going to saturation (see plot).

TCS
at 13:00UTC we lost the data from the TCS chiller guardian (expert informed).

Images attached to this report
Search Help
×

Warning

×